HUBBLE SPACE TELESCOPE DAILY REPORT #2133 PERIOD COVERED: 0000Z (UTC) 05/28/98 - 0000Z (UTC) 05/29/98 Daily Status Report as of 149/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled with no anomalies. 1.3 Completed Six Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.4 Completed NIC/2 7495 (Survey of Gravitational Lenses as Cosmological Tools) The Near Infrared Camera (Camera Two) was used to make survey observations of object HST14176+5 by obtaining deep, high-resolution images in the optical and near IR in order to see if it is gravitational lensed and could be used as a cosmological tool. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.6 Completed Three Sets of WF/PC-2 7336 (Cosmological Parameters Omega and Lambda From High-Redshift Type Ia Supernovae) The WF/PC-2 was used to make observations of the high redshifted Type Ia supernovae 14H0.5Z2, 11H0.5Z2 and 14H0.5Z1. HST can make a direct measurement of the cosmological parameters, Omega, Lambda, and curvature, using Type Ia supernovae as calibrated "standard candles." The observations were executed as scheduled with no anomalies. 1.7 Completed Four Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. The observations were completed as planned, and no anomalies were reported. 1.8 Completed Three Sets of NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-5135, NGC-4314 and NGC-6221 far from the center by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The proposal completed as expected with no anomalies. 1.9 Completed STIS/MA1 7308 (HST FAR-UV Imaging and Spectra of Jupiter's Aurora Coordinated With Galileo) The WF/PC-2 and the Space Telescope Imaging Spectrograph (MA2) were used to make far-ultraviolet imaging and spectral observations of the aurora near both poles of the planet Jupiter at the same time as scheduled GALILEO spacecraft measurements of Jupiter's magnetospheric properties and ultraviolet emissions over summer/fall 1997. The observations were completed as planned with no anomalies. 1.10 Completed STIS/CCD 7911 (STIS Galactic Plane Pure Parallel Program) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude survey observations in parallel with the prime science observations. Short exposures with the 50CCD imaging mode will provide the wavelength zero-point calibration. These data will be placed immediately into the Hubble Data Archive. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed WF/PC-2/NIC/1/2/STIS/CCD/MA2 7560 (STIS GTO Observations) The WF/PC-2, the Near Infrared Camera (Cameras One and Two) and the Space Telescope Imaging Spectrograph (CCD and MA2) were used to make observations of galaxy NGC-3396. The solar blind and long slit capabilities of STIS will be used to investigate the population of hot stars in galaxies. The observations were executed as scheduled with no anomalies. 1.12 Completed WF/PC-2 7282 (A Continuing Snapshot Survey for Compact Stellar Groups and Clusters around Wolf-Rayet Stars) The WF/PC-2 was used to make observations of the Wolf-Rayet star WR137 as part of a continuing snapshot survey of compact stellar groups and clusters. The newly imaged clusters allow us to accurately determine the distances, luminosities, and ages of the WR members. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed NIC/3 7818 (Spectrophotometry of Pluto and Charon) The Near Infrared Camera (Camera Three) was used to make high signal-to-noise ratio specrophotometric observations of the planet Pluto and its moon Charon. These observations will be used to separate the spectroscopic signatures of each object in future ground-based combined light measurements. These spectra will also provide a very sensitive search for trace molecular components on the surfaces of Pluto and Charon. The observations were completed as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 14 Successful : 14 Scheduled Reacquisitions : 4 Successful : 4 2.2 FHST Updates: Scheduled : 32 Successful : 30 Per HSTAR 6496, the FHST roll delay updates at 149/0228Z and at 149/0351Z failed. 2.3 Operations Notes: Using ROP SR-1, the EDAC error counter was cleared at 148/1414Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH